OpenStudy (anonymous):

Medal to the best answer!! How many solutions are there to the following system of equations? 2x + 6y = –12 10x + 32y = –62 A. 0 B. infinitely many C. 1 D. 2

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

144

OpenStudy (anonymous):

@OOOPS

OpenStudy (anonymous):

I think Mr.Mathmale would Explain better :)

OpenStudy (mathmale):

The question here boils down to: These two lines intersect 1) nowhere 2) in one point 3) at infinitely many points You could answer this question by graphing the two lines, or you could solve the system of equations algebraically: elimination by substitution, elimination by addition/subtraction. Your choice.

OpenStudy (anonymous):

I think it's B :) :(

OpenStudy (wolf1728):

A) 2x + 6y = –12 B) 10x + 32y = –62 Multiply equation A by -5 to get -10x -30y = 60 Then add this to B)

OpenStudy (wolf1728):

The answer isn't B infinitely many

OpenStudy (anonymous):

Oh

OpenStudy (wolf1728):

Look at my previous reply Equation A becomes A) -10x -30y = 60 then add this to B) 10x + 32y = –62

OpenStudy (anonymous):

A?

OpenStudy (wolf1728):

Instead of just guessing letters it would be easier just to solve the equations.

OpenStudy (anonymous):

D?

OpenStudy (wolf1728):

I think you just might be guessing letters again. (Hint - the answer is NOT Z either)

OpenStudy (anonymous):

It's Either C or B

OpenStudy (mathmale):

To be honest, I'm sorry and disappointed to see that you're trying to guess the answer. It'd be so much more interesting, and so much more helpful, to learn how to do these calculations. Most of us on OpenStudy won't tell you the answer or confirm a guess for a reason: we'd like for you to develop the skills to solve these problems.

OpenStudy (anonymous):

oh

OpenStudy (mathmale):

Have you tried graphing the two equations? Have you tried finding the slopes of the two lines and comparing them?

OpenStudy (wolf1728):

Okay let's solve the equations: A) -10x -30y = 60 B) 10x + 32y = –62 that sums to 2y = -2 So y = -1 Putting that into 2x + 6y = –12 we get 2x -6 = -12 2x = -6 x = -3 So we have 2 solutions to these equations and the answer is D 2

OpenStudy (anonymous):

Thanks

OpenStudy (wolf1728):

u r welcome